OpenStudy (anonymous):

A sample of a carbon-hydrogen-oxygen compound is 12.7% hydrogen and 36.2% oxygen. how many grams of carbon are there in a 12.2 g sample of the compound?

OpenStudy (anonymous):

%C = 100 - %H - %O = 100 - 12.7 - 36.2 = 51.1 Then mC = sample x %C = 12.2 x 51.1% = 6.23 (g)